Restaurant “King Steak” offers a set lunch of a 150 grams steak for $100. The steaks are cut by a machine
and the weight of a steak can be assumed to be normally distributed with mean 152 grams and standard
deviation 1.7 grams.
(a) What is the probability that a customer gets a steak that weighs less than 150 grams?
(b) There is a 4% chance that a customer gets a steak that weighs less than W grams. What is the value of
W?
(c) There is a table with 8 customers, everyone orders this set lunch. What is the probability that there are
exactly 2 customers whose steaks weigh less than 150 grams?
(d) Sometimes there are customers ordering two sets of lunch. Use T to denote the total weight of two
steaks. Find the mean, variance, and standard deviation of T.
1
Expert's answer
2020-04-02T14:29:46-0400

Let "X=" the weight of a steak: "X\sim (N, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"

Given that "\mu=152\ g, \sigma=1.7\ g"

(a) What is the probability that a customer gets a steak that weighs less than 150 grams? 


"P(X<150)=P(Z<{150-152\over 1.7})\approx"

"\approx P(Z<-1.17647)\approx0.1197"

(b) There is a 4% chance that a customer gets a steak that weighs less than W grams. What is the value of W? 


"P(X<W)=P(Z<{W-152\over 1.7})=0.04"

"{W-152\over 1.7}\approx-1.7507"

"W\approx149\ g"

(c) There is a table with 8 customers, everyone orders this set lunch. What is the probability that there are exactly 2 customers whose steaks weigh less than 150 grams? 

Let "Y=" the number of customers whose steaks weigh less than 150 grams: "Y\sim Bin (n,p)"


"P(Y=y)=\binom{n}{y}p^y(1-p)^{n-y}"

Given that "n=8, p=0.1197"

"P(Y=2)=\binom{8}{2}(0.1197)^2(1-0.1197)^{8-2}\approx0.1867"

(d) Sometimes there are customers ordering two sets of lunch. Use T to denote the total weight of two

steaks. Find the mean, variance, and standard deviation of T.


"T\sim N(\mu+\mu,\sigma^2+\sigma^2)"

Then

"\mu_T=150\ g+150\ g=300\ g"

"Var(T)=\sigma_T^2=1.7^2+1.7^2=5.78"

"\sigma_T=\sqrt{5.78}\ g=1.7\sqrt{2}\ g\approx2.4\ g"